OpenStudy (anonymous):

Part 1: Create your own quadratic equation that cannot be solved by factoring, but can be solved using the quadratic formula. Identify the values of a, b, and c, and find the solutions using the quadratic formula. Show all work to receive credit. Part 2: Using complete sentences, explain how you know that the equation from Part 1 cannot be solved by factoring, but can be solved by using the quadratic formula.

OpenStudy (anonymous):

sorry I meant x^2+4x^2+5x^2=0

OpenStudy (the_fizicx99):

... That's 10x^2 Hmm, how about: x^2 + 4x + 7 = 0 there solve!

OpenStudy (anonymous):

lol it was my mistake putting the ^2 to the one of the middle and the last one sorry

OpenStudy (anonymous):

x^2 + 4x + 7 = 0 a: 1 b: 4 c: 7 x = -4 + sqrt4^2 - 4(1)(7)/ 2(1) x = -4 + sqrt16 - 4(1)(7)/ 2(1) x = -4 + sqrt16 - 28/ 2(1) x = -4 + sqrt16 - 28/ 2 x = -4 + sqrt-12/ 2 x= -4 + sqrt-12/2 x= -4 - sqrt -12/2 x= -4 + -3.46/2 x= -4 -3.46/2 x= -4 + -3.46/2 x= -7.46/2 Help with the last simplifications"? @tHe_FiZiCx99